OpenStudy (anonymous):

The graph of which equation has a vertex of (–2, 0)?

OpenStudy (anonymous):

there are many graphs that has (-2, 0) as it's vertex... do you have choices?

OpenStudy (anonymous):

Answer y = 3(x + 2)2 y = 3(x + 2)2 + 1 y = (x – 2)2 + 2 y = 3(x – 2)2

OpenStudy (precal):

 (h,k) is your vertex